The data below show the wages earned by six workers at three different firms (A, B, and C). The firms name is shown in column 1 and the workers wages in column 2. The third column shows the number of times that each firm has been audited for financial fraud while the worker was employed at the firm. The fourth column is a binary variable indicating whether or not the employee works at the accounting department. In the following questions you may check your work using a calculator but must show how you set up the expression you are solving (i.e. show your work). 1. firm wage audits accountant 14 19 15 12 16
3. Properties of scaling. Suppose that both wages and audits are three times larger for each individual in the data above. a) What is the new mean and variance of wages? How do these compare to your answer in b) What is the new covariance of wages and audits? c) Prove mathematically that the Cov(ax, aY) - aCov(x.Y) using the properties of part a) of Question 1)? Be precise. in part d) of Question 1)? summation. Show each step. How does this compare to your answer
Answer #1

New data when the wages and audit is larger 3 times is

firm wage audits
A 33 9
A 42 6
B 57 0
B 45 0
C 36 3
C 48 6

Using the formula, \text{mean of wages}=\mu={\sum_{i=1}^{6} x_i\over 6}=43.5 and  

Also \text{Variance in wages}=\sigma^2={\sum_{i=1}^{6}(\mu- x_i)\over 6}= 62.25

The earlier mean was 14.5, and new mean is 43.5, clearly 3 × 14.5-43.5

and earlier varaince is 6.916667 and new varaince is 62.25, clearly 32 × 6.916667= 62.25

b) The new covaraince is given by

\text{Cov}=\sum_{i=1}^{6}(x-\mu)(y-\mu_1)

Where y denotes audit, and \mu_1 its mean.

Using the new values we have new Cov-_16.5

Earlier covariance is -1.833333333, and 3^2\times -1.8333333= -16.5

c) Let X=x_1,x_2,\cdots, x_n and  Y=y_1,y_2,\cdots, y_n then

aX=ax_1, ax_2,\cdots, ax_n~~~,~~~aY=ay_1,ay_2,\cdots,a y_n

Then mean of X is \mu and mean of Y is \mu_1 . Also mean of aX is a\mu and mean of aY is a\mu_1

\text{Cov}(aX,aY)=\sum_{i=1}^{n}(ax_i-a\mu)(ay_i-a\mu_1) \\\\ ~~~~~~~~~~ ~~~~~~~~~~~~=\sum_{i=1}^{n}a^2(x_i-\mu)(y_i-\mu_1)\\\\ ~~~~~~~~~~ ~~~~~~~~~~~~=a^2\sum_{i=1}^{n}(x_i-\mu)(y_i-\mu_1)\\\\ ~~~~~~~~~~ ~~~~~~~~~~~~=a^2\text{Cov}(X,Y)